OpenStudy (littleporker):

For y = 4x + 2x2 – x3, show how to find the value for y, when x = 2. Show your work

OpenStudy (wolf1728):

In that equation, I'd say the 2x2 means "2x squared" and x3 means "x cubed" correct?

OpenStudy (littleporker):

Yes that's right! Sorry I didn't put that!

OpenStudy (wolf1728):

We have the equation y = 4x + 2x2 – x3 and we want to determine "y" when x=2 4*2 =? + 2 * 2 * 2 = ? -2 * 2 * 2 =?

OpenStudy (wolf1728):

Okay what is 4*2 ?

OpenStudy (littleporker):

Sorry I've been busy! And I actually worked out the problem and got 8!! Is that correct??

TheSmartOne (thesmartone):

Correct :)

OpenStudy (littleporker):

Thank You!!!!

TheSmartOne (thesmartone):

Anytime! (:

OpenStudy (littleporker):

Hey maybe you can help me? I don't know how to ask another question on here? Can you tell me how? Thanks!!!!

TheSmartOne (thesmartone):

You'll need to "close" the current question to open another one :)

OpenStudy (littleporker):

Okay thanks again!!!!